The total number of Rs.5 and Rs10 coins in my bag is 20.If i have Rs 145 in my bag.Lets calculate the number of coins of each kind i had in my bag.
Answer by CubeyThePenguin(3113)

N = number of Rs. 5 coins
D = number of Rs. 10 coins

N + D = 20
5N + 10D = 145

Multiply the first equation by 10.

10N + 10D = 200
5N + 10D = 145

Subtract to get 5N = 55, so N = 11.

There are 11 five-Rs. coins and 9 ten-Rs. coins.